# MAPO: Advancing Multilingual Reasoning through Multilingual Alignment-as-Preference Optimization

### 12 Jan 2024 v1

@@ 5 Experiment Results
5.1 Preference Optimization Advances Multi-lingual Reasoning Ability Effectively
Based on the results in both three Table 1,2,3, our method has effectively enhanced the reasoning capabilities of the base model. Across the average non-English reasoning accuracy rates on three datasets, our approach has outperformed ChatGPT and achieved state-of-the-art performance. Specifically, even on the test set MGSM, which corresponds to the base model fine-tuning data, our method still effectively improved the reasoning accuracy by 3%. This enhancement effect is even more significant on the out-of-domain dataset MSVAMP, where we achieved an average accuracy improvement of 16.6%. Meanwhile, MultiLingual-RFT, using the same sampling results, only get an improvement of 8.7%. This demonstrates the efficiency of our method.
From the perspective of individual languages, all languages have experienced an increase in capabilities through alignment. The effect is relatively significant on some low-resource languages that previously underperformed. Our method has increased the accuracy on MSVAMP for Bengali, Thai, Swahili, and Japanese by 21.1%, 19.3%, 16.6%, and 18.7%, respectively. Moreover, the medium-to-high resource languages, which originally performed relatively well on the base model, have also continuously gained improvements in capabilities as the alignment progresses.
It’s surprising that our training data does not contain English, but the accuracy on English questions has been greatly improved. We believe this is primarily due to the alignment facilitating a more consistent understanding of reasoning across different languages, which contributes to the enhancement of its capabilities, including performance in English. Besides, our method also achieved considerable improvements on the MetaMathOctopusP 7B, which has stronger English reasoning capabilities, proving the robustness of our approach across different base models.
@@ 5.2 Alignment is the Key to Enhanced Multilingual Reasoning
Regarding the alignment of the reasoning process, as shown in Figure 3, our method effectively improves the consistency of the reasoning process, particularly for languages where the base model and ChatGPT had poorer alignment, such as Bengali, Thai, and Japanese. This proves that the model can show more similar reasoning thinking to English on non-English after alignment.
Additionally, we notice that alignment also contributes to more consistent final answers. A higher ARC indicates that there is a greater overlap between the questions answered correctly in non-English and those answered correctly in English. From the results in Figure 3, our alignment framework has greatly increased the ARC for each language. This means that the performance gains in Table 1 stem from the parts that intersect with English. These observations demonstrate that our method effectively aligns the model’s non-English reasoning processes with English, thereby enhancing reasoning abilities in non-English languages.
5.3 Preference Optimization Brings a More Generalizable Improvement in Multilingual Reasoning Ability
As indicated in Table 3, both SFT and DPO have achieved performance improvements on the test set of the MNumGLUESub. However, on the out-of-domain test set MSVAMP, DPO achieved a 13.7% increase in accuracy, which is significantly higher than the 7.9% increase achieved by SFT. More surprisingly, while DPO effectively improving the multilingual reasoning performance on both two datasets, it also maintained competitive reasoning performance on MGSM. Iterative DPO could even lead to further enhancements in performance. In contrast, SFT exhibited a 4.7% performance degradation.
We believe that this is mainly because model directly learns the given labels during SFT. The labels inevitably have distributional characteristics unique to the dataset, which are difficult to generalize to other datasets. Conversely, our method does not give the model "correct answers", but rather guides the model to generate outputs that are more aligned with the alignment characteristics in the space based on preferences, which is more effective in allowing the model to master the multi-language alignment learning objective, and this preference can better generalize to out-of-domain data sets.
@@ 6 Analysis
GSM-INS MNUM-Sub
Valid Sample 575 895
En Self-Bleu 0.81 0.53
Non-En Self-Bleu 0.84 0.60
Avg Acc on MSVAMP 31.6 54.4
Table 4: Comparison between GSM8KINSTRUCT (GSM-INS) and MNumGLUESub(MNUM-Sub)
@@
6.1 Why Not Sample on SFT Dataset
A natural question is whether we can simply sample and optimize preference on the SFT dataset (like GSM8KINSTRUCT). Experimental results indicate that the SFT dataset is not suitable as a sampling dataset for preference optimization. We randomly choose 1000 samples from the sampling results of both the GSM8KINSTRUCT and MNumGLUESub datasets. To ensure fairness, we define the samples with more than 5 answers after deduplication as valid sample and calculated the self-bleu Zhu et al. (2018) on first 5 answers within the samples (self-bleu calculates the overlap between a group of text contents, which is widely used to measure the diversity of generated results).
The analysis results are shown in the Table 4. The valid sample collected on the SFT dataset is significantly reduced, and the self-bleu is significantly higher than that of other datasets. This effectively proves that SFT will make the model quickly remember the given standard answer, which leads to a serious degradation in the diversity of the generation. The policy model struggles to learn the alignment preferences we truly hope to optimize and it easily falls into the trap of overfitting. Fortunately, compared to complex reasoning processes, multilingual questions and instructions are relatively easy to obtain, and their construction costs are lower. We only use MNumGLUE which has only 1700 samples, and the prospect of sampling based on more multilingual questions and optimizing for alignment preferences remains for future research.
@@ 6.1 Why Not Sample on SFT Dataset
Figure 4: Average accuracy on Non-English math problems and ACR versus preference optimization step on MSVAMP. MathOctopusP 13B and ChatGPT are selected for comparison
6.2 LLM Gradually Learn to Align and Improve Multilingual Reasoning during Preference Optimization
To better analyze the state of the model during the optimization preference process, we tested the checkpoints of the DPO and PPO-LoRA training on MSVAMP and visualized the accuracy and ACR respectively in Figure 4.
The results show that the optimization of DPO is quite efficient. DPO enabled the MathOctopusP 7B model to achieve a higher accuracy than both ChatGPT and MathOctopusP 13B within just 100 steps. As the optimization continues, the accuracy remains relatively stable while the degree of alignment still shows an increasing trend.
On the other hand, although the optimization speed of PPO is slower, it took fewer than 1000 steps to surpass the larger-scale MathOctopusP 13B. As the model continued to explore and optimize preferences, it achieved superior multilingual reasoning performance compared to ChatGPT at 1600 steps.
The experimental results demonstrate that the model, through exploration and preference optimization, continuously aligns non-English reasoning with English reasoning, effectively enhancing its multilingual reasoning capabilities. For PPO-LoRA, although optimization was halted around 2700 steps due to computational limitations, it is observable that the growth in multilingual reasoning consistency and ability continues even beyond 2500 steps, showcasing the potential of our preference optimization framework.
6.3 Case Study
In the table example, we present an instance illustrating the impact of alignment on the model’s output. The baseline model, when presented with a question in English, correctly determined that an addition operation was required to solve for the suggested number of students to include bacon. However, when faced with the same question expressed in Chinese, it incorrectly assessed the variable computation relationship. This further demonstrates that although the original model was fine-tuned for multiple languages, there is an inconsistency in its internal reasoning capabilities and thought processes, leading to errors in non-English reasoning. After alignment, the model’s overall reasoning process became more similar to its English responses, and it correctly identified the numerical relationship. This example directly illustrates how our framework improves reasoning abilities by aligning the reasoning in other languages with English.
@@ 6.3 Case Study
Figure 5: Accuracy across ten languages on MSVAMP after training MathOctopusP7B on preference datasets constructed using translation models of different scales.
6.4 Robustness over Different Scale of Translation Model
@@
We also conducted ablation experiments comparing the translation models used during the preference estimation phase. In addition to NLLB-600M-distilled, we employed NLLB-1.3B-distilled  [^5] https://huggingface.co/facebook/nllb-200-distilled-1.3B and NLLB-3.3B  [^6] https://huggingface.co/facebook/nllb-200-3.3B, keeping the sampling dataset the same, and used them to estimate preferences, repeating the aforementioned experiments. The results are shown in Figure 5, where we found that all three models achieved impressive results, showing better reasoning capabilities over the Base Model and ChatGPT across ten languages. These results demonstrate that our method is robust for translation models of various sizes.
7 Conclustion
@@
In this paper, we propose a novel multilingual alignment-as-preference optimization framework, significantly enhancing reasoning ability in non-pivot languages by aligning them with pivot languages. Specifically, we utilize an open-source translation model to estimate alignment preferences, which are then optimized using either DPO or PPO. Experimental results demonstrate that our framework achieved improvements in reasoning accuracy across three test sets, with a notable 13.7% increase in reasoning accuracy on out-of-domain datasets. The analysis also confirms that enhancing alignment through our method is the key to improvements in multilingual reasoning capabilities. Furthermore, we found that iterative DPO contributes to further alignment and enhancement of reasoning abilities, pushing the improvement to 16.7%.
Appendix A Appendix
A.1 Experiment Details
@@
Our code is primarily based on the trl 777 https://github.com/huggingface/trl , with some minor modifications made. The modified code will also be made available at  [^8] https://github.com/NJUNLP/MAPO
@@
LoRA: For experiments using LoRA, such as DPO LoRA and PPO LoRA, we standardize the experimental settings, optimizing the [q_proj, v_proj , o_proj] modules. The values of r and alpha are set to 128 and 256, respectively.
DPO: We employ a learning rate of 1e-6, with $\beta$ set at 0.1, and a warmup step count of 100. The batch size is configured to 128. The optimization process is capped at a maximum of 1000 steps, and we save the checkpoint that corresponds to the lowest loss on validation set.
PPO: We have configured a learning rate of 2e-5, with a batch size of 64 and 2 PPO epochs. We utilize the AdamW optimizer to improve the stability of the optimization, setting the epsilon value to 1e-5. Additionally, we have implemented a linear learning rate warm-up technique for the first 150 steps. All other hyperparameters are set to the default values provided by the trl library. This is an appendix.

### 22 Feb 2024 v2

@@ 5 Experiment Results
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
GPT-3.5-Turbo 31.2 38.0 40.0 36.0 44.0 43.2 46.0 47.2 41.6 54.4 42.2
MathOctopus 13B 42.4 39.2 44.8 38.8 49.6 45.2 48.4 53.6 43.2 54.8 46.0
MAmmoTH ${}^{\dagger}$ 3.6 4.8 2.4 10.8 17.2 26.0 33.2 32.4 32.8 49.6 21.3
WizardMath ${}^{\dagger}$ 2.0 4.0 3.4 24.0 22.4 30.8 30.4 34.8 30.4 47.6 23.0
MetaMath 7B 6.4 4.0 3.2 39.2 38.8 47.2 56.8 58.0 52.8 63.2 37.0
MathOctopus 7B 29.2 33.6 36.4 35.2 39.2 38.8 44.8 42.4 43.2 52.0 39.5
+ m-RFT 25.6 31.2 28.8 34.0 39.2 36.0 34.8 34.4 36.4 43.2 34.4
+ MAPO-DPO(ours) 30.8 38.0 37.6 45.2 47.2 42.0 45.2 43.2 40.8 45.6 41.6
MetaMathOctopus 7B 25.6 42.8 36.4 40.0 46.4 46.8 49.6 54.4 46.4 66.4 45.5
+ m-RFT 23.2 33.6 34.0 34.0 47.2 43.2 45.6 47.6 44.8 62.8 41.6
+ MAPO-DPO(ours) 36.0 44.8 44.8 47.6 55.2 53.6 53.6 56.8 52.4 70.8 51.6
Table 3: Model Performances on MGSM test set. “Avg” represents the average performance in ten languages and bold text denotes the best results within the same model size. Results marked with ${}^{\dagger}$ come from Chen et al., 2023.
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
GPT-3.5-Turbo 36.2 42.6 47.2 58.1 60.6 42.6 41.5 54.9 39.4 70.6 49.4
MathOctopus 13B 32.2 35.8 38.2 42.7 46.0 38.8 36.0 45.2 37.5 50.5 40.3
MAmmoTH 7B 1.7 4.7 4.5 23.5 29.9 26.9 34.7 37.9 36.5 42 24.2
WizardMath 7B 9.8 13 6.2 31.8 36 34.3 34.1 39.9 36.5 45.4 28.7
MetaMath 7B 13.2 19.2 12.1 50.7 53.3 52.9 54.2 56.5 56.7 63.3 43.2
MathOctopus 7B 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 46.9 37.1
+ m-RFT 38.0 42.9 41.8 48.2 51.6 45.2 42.9 49.3 42.9 51.2 45.4
+ MAPO-DPO(ours) 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 55.9 50.4
MetaMathOctopus 7B 24.5 35.2 36.2 44.1 45.8 39.2 32.6 45.2 36.7 52.5 39.2
+ m-RFT 40.3 45.0 45.2 51.4 57.8 51.6 51.6 58.8 50.1 65.3 51.7
+ MAPO-DPO(ours) 40.7 46.0 45.0 58.2 59.3 53.1 51.4 57.4 52.0 66.1 52.9
Table 4: Model Performances on MNumGLUESub test set. “Avg” represents the average performance in ten languages and bold text denotes the best results within the same model size.
5.1 Preference Optimization Improves Multilingual Reasoning Effectively
Experimental results in Table 2,3,4 consistently demonstrate that our method has effectively enhanced the reasoning capabilities of the base model and achieved state-of-the-art performance. More specifically, the improvement is most impressive on the out-of-domain dataset MSVAMP, where we achieved an average accuracy improvement of 16.2% and 11.7%, over the two base models, MathOctopus7B and MetaMathOctopus7B, respectively. We also observe significant improvements in MGSM and MNumGLUESub.
From the perspective of languages, most languages have improvement after alignment and it is more significant in some low-resource languages that previously underperformed. Our method has increased the accuracy on MSVAMP for Bengali, Thai, Swahili, and Japanese by 21.1%, 19.3%, 16.6%, and 18.7%, respectively.
Surprisingly, our preference optimization dataset does not contain English, but the accuracy of English has also improved. After alignment, MetaMathOctopus even surpassed the English mathematical reasoning models on English questions across three datasets. We believe this is primarily due to the alignment facilitating a more consistent understanding of reasoning across different languages, which contributes to the enhancement of general reasoning capabilities.
Meanwhile, our approach has also achieved stable and impressive improvements on the MetaMathOctopus 7B which has stronger reasoning capabilities, propelling it past the ChatGPT and the larger scale model MetaMathOctopus13B, demonstrating the robustness and potential of our framework.
@@ 5.2 Alignment is the Key to Enhanced Multilingual Reasoning
Regarding the alignment of the reasoning process, as shown in Figure 3, our method effectively improves the consistency of the reasoning process, particularly for languages where the base model had poorer alignment, such as Bengali, Thai, and Japanese. This proves that the model can show more similar reasoning thinking to English on non-English after alignment.
Additionally, we notice that alignment also contributes to more consistent final answers. A higher ACR indicates that there is a greater overlap between the questions answered correctly in non-English and those answered correctly in English. From the results in Figure 3, our alignment framework has greatly increased the ACR for each language. This means that the performance gains in Table 2 stem from the parts that intersect with English. These observations demonstrate that our method effectively aligns the model’s non-English reasoning processes with English, thereby enhancing reasoning abilities in non-English languages.
5.3 Generalizable Multilingual Reasoning
As indicated in Table 2, on the out-of-domain test set, MAPO achieved a 16.2% and 11.7% increase in accuracy, which is significantly higher than the 7.5% and 3.7% achieved by m-RFT. Surprisingly, while MAPO effectively improves the multilingual reasoning performance on both two datasets, it also improves the performance on MGSM. In contrast, m-RFT exhibited a 5.1% and 3.9% performance degradation.
We believe that this is mainly because the model directly learns the given labels during SFT. The labels inevitably involve data-specific attributes, which are difficult to generalize. Conversely, our method does not give the model “correct answers”, but instead guides the model to generate outputs we prefer, which is more effective in allowing the model to learn multilingual reasoning, and this preference can better generalize across various challenges.
@@ 6 Analysis
MGSM MNumGLUESub
Acc. on MSVAMP 31.6 54.9
Self-BLEU(En) 0.81 0.53
Self-BLEU(Non-En) 0.84 0.60
Table 5: Comparison between MGSM8KInstruct (MGSM) and MNumGLUESub. We report the average Self-BLEU scores for the sampled reasoning in English and non-English.
@@
6.1 Preference Estimating on a Dataset Different from SFT
We conduct experiment with preference optimization using the MGSM8KInstruct dataset instead of the MNumGLUESub. Table 5 present the average accuracy over 10 languages on MSVAMP. The results indicate that preference optimization based on the SFT dataset lead to the performance degradation on the out-of-domain test set MSVAMP. We believe this is because SFT make the model quickly remember the oracle reasoning process, which leads to a serious degradation in the generation diversity. Thus, the policy model struggles to learn the alignment preferences we truly hope to optimize and easily falls into the trap of overfitting.
For further analysis, We randomly sample 1000 reasoning processes of LLMs on the MGSM8KInstruct and MNumGLUESub, and calculated the Self-BLEU Zhu et al. (2018) on sampled reasoning process of the same problem. The results (Table 5) suggest that on MGSM dataset, LLMs tends to generate quite similar reasoning, compared with MNumGLUESub, which confirms the diversity degradation.
@@ 6.1 Preference Estimating on a Dataset Different from SFT
Figure 4: Average accuracy on Non-English math problems and ACR versus preference optimization step on MSVAMP. MathOctopus 13B and ChatGPT are selected for comparison.
@@
Figure 5: Accuracy across ten languages on MSVAMP after training MathOctopus 7B on preference datasets constructed using translation models of different scales.
6.2 Scaling of Translation Model
@@
We also conducted ablation experiments comparing the translation models used during the preference estimation phase. In addition to NLLB-600M-distilled, we employed NLLB-1.3B-distilled and NLLB-3.3B to estimate the preference and repeat the experiments. The results are shown in Figure 5, where we find that all three models achieved impressive results, showing stable improvements over the Base Model and ChatGPT across ten languages. These results demonstrate that our method is robust for translation models of various sizes.
6.3 Reasoning alignment during Preference Optimization
To better analyze the optimization preference process, we tested the checkpoints of the DPO and PPO-LoRA on MSVAMP and visualized the accuracy and ACR respectively in Figure 4.
The results show that the optimization of DPO is quite efficient. DPO enabled the MathOctopus 7B model to achieve higher accuracy and ACR than both ChatGPT and MathOctopus 13B within just 100 steps. As the optimization continues, the accuracy remains relatively stable while the degree of alignment still shows an increasing trend.
Meanwhile, although the optimization speed of PPO is slower, it take fewer than 1000 steps to surpass MathOctopus 13B. As the model continued to explore and optimize the preferences, it achieved superior multilingual reasoning performance compared to ChatGPT at 1600 steps.
The experimental results demonstrate that the model continuously aligns non-English reasoning with English reasoning, effectively enhancing its multilingual reasoning capabilities. For PPO-LoRA, although optimization is halted around 2500 steps due to the limitations on computational resources, we notice that growth in multilingual reasoning consistency and ability continues even beyond 2500 steps, indicating the potential of our preference optimization framework.
6.4 Alignment Reasoning for Improvements
The example in Table 6 directly illustrates how our framework can improve reasoning ability by aligning reasoning in other languages to English. For the given English problem, the basic model correctly analyzes that an addition should be conducted to find the number of students recommended to join Bacon. However, when given the same problem in Chinese, LLM misjudges the relationship of variables and writes a significantly different reasoning process. This directly proves that although the baseline model has been fine-tuned with multilingual reasoning data, its reasoning and thoughts are inconsistent, whereas non-English reasoning is more prone to errors. After alignment, the reasoning thoughts are more similar to the English answer, and the reasoning thoughts are also corrected.
7 Conclusion
@@
In this paper, we propose MAPO, a novel multilingual alignment-as-preference optimization framework, enhancing reasoning ability in non-dominant languages by aligning them with dominant languages. Experimental results demonstrate that our framework achieves significant improvements on various base models across all three benchmarks, especially with a notable 16.2% increase in average accuracy on the out-of-domain datasets MSVAMP. The analysis confirms that enhancing alignment through our method is the key to improvements in multilingual reasoning capabilities.
Limitation
Similar to previous work on preference optimization, our method necessitates a policy model that has the preliminary multilingual reasoning capabilities through SFT. Meanwhile, due to limitations in computational resources, our experiments are confined to aligning models of 7B size and exploring two preference optimization algorithms, PPO and DPO. Should resources permit, we aim to extend our exploration to models of 13B and 70B sizes and examine the performance of a broader spectrum of preference optimization algorithms.
Appendix A Experiment Details
@@
Our code is primarily based on the trl 666 https://github.com/huggingface/trl , with some minor modifications made. The modified code will also be made available at our project.
@@
LoRA: For experiments using LoRA, such as PPO LoRA, we optimize the [q_proj, v_proj , o_proj] modules. The values of r and alpha are set to 128 and 256, respectively.
DPO: We employ a learning rate of 1e-6, with $\beta$ set at 0.1, and a warmup step count of 100. The batch size is configured to 128. The optimization process is capped at a maximum of 1000 steps, and we save the checkpoint that corresponds to the lowest loss on validation set. The training takes around 4 hours on 8 H100 GPUs.
PPO: We have configured a learning rate of 2e-5, with a batch size of 64 and 2 ppo epochs. We utilize the AdamW optimizer to improve the stability of the optimization, setting the epsilon value to 1e-5. Additionally, we have implemented a linear learning rate warm-up technique for the first 150 steps. All other hyper parameters are set to the default values provided by the trl library. We set the maximum optimization steps to 2600 and report the results at this checkpoint.
Appendix B Supplemental Experiment Results
To verify the robustness of our method, we conducted experiments on different preference optimization algorithms. Due to the limited computational resources, we optimized the PPO algorithm using LoRA. The detailed experimental results of PPO LoRA and each round in Iterative DPO are shown in Table 7.
Experiments demonstrate that our framework achieve an effective improvement in multilingual reasoning capabilities based on both PPO and DPO. Despite the limited computational resource, where PPO only updated a part of parameters with merely 2600 steps, it has already brought impressive performance enhancements on all three datasets. Additionally, with the increasing rounds of DPO, the model exhibited progressively stronger multilingual reasoning abilities that has not yet reached its limit, revealing the potential of our approach.
Appendix C Used Scientific Artifacts
Below lists scientific artifacts that are used in our work. For the sake of ethic, our use of these artifacts is consistent with their intended use.
  • [itemsep=1pt]
  • Transformers (Apache-2.0 license), a framework to facilitate downloading and training state-of-the-art pretrained models.
  • trl (Apache-2.0 license), a full stack library where provide a set of tools to train transformer language models with Reinforcement Learning. The library is built on top of the Transformers library.
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
Benchmark: MSVAMP
MathOctopus 7B 27.7 35.9 39.4 41.6 42.7 44.2 44.0 45.1 45.3 46.4 41.2
+ PPO LoRA 38.9 47.5 47.1 51.0 51.7 51.1 50.3 51.6 51.4 52.7 49.3
+ DPO Iter1 42.0 53.2 52.7 54.7 56.4 56.9 55.7 58.5 59.3 59.6 54.9
+ DPO Iter2 45.3 53.9 53.8 56.8 58.1 56.6 58.7 59.1 58.5 60.0 56.1
+ DPO Iter3 48.8 55.2 56.0 60.3 58.8 58.3 58.1 59.7 60.8 58.4 57.4
MetaMathOctopus 7B 36.1 47.5 49.4 51.3 54.5 53.6 56.6 60.0 57.2 64.2 53.0
+ PPO LoRA 45.7 52.0 52.2 61.2 58.3 57.0 58.7 60.7 61.7 67.5 57.5
+ DPO Iter1 44.1 58.2 59.0 60.3 62.3 63.5 65.1 63.7 64.2 70.2 61.1
+ DPO Iter2 48.5 61.8 59.2 64.3 64.3 64.4 65.0 66.1 65.4 70.9 63.0
+ DPO Iter3 50.1 61.6 61.7 65.9 65.7 64.8 68.4 68.5 68.6 71.6 64.7
Benchmark: MGSM
MathOctopus 7B 29.2 33.6 36.4 35.2 39.2 38.8 44.8 42.4 43.2 52.0 39.5
+ PPO LoRA 31.2 38.4 38.4 37.2 43.6 35.2 46.0 44.0 38.8 51.2 40.4
+ DPO Iter1 29.2 36.4 35.6 35.6 41.6 38.4 40.8 42.0 37.6 46.8 38.4
+ DPO Iter2 30.4 36.0 37.6 38.0 45.2 39.6 42.0 47.6 41.2 45.2 40.3
+ DPO Iter3 30.8 38.0 37.6 45.2 47.2 42.0 45.2 43.2 40.8 45.6 41.6
MetaMathOctopus 7B 25.6 42.8 36.4 40.0 46.4 46.8 49.6 54.4 46.4 66.4 45.5
+ PPO LoRA 36.0 41.2 41.6 46.4 54.8 53.6 54.0 55.6 51.6 68.0 50.3
+ DPO Iter1 32.8 43.2 40.4 48.8 49.2 52.8 54.4 52.8 50.0 64.8 48.9
+ DPO Iter2 34.0 48.0 45.2 40.4 54.0 52.0 50.8 54.0 49.2 70.4 49.8
+ DPO Iter3 36.0 44.8 44.8 47.6 55.2 53.6 53.6 56.8 52.4 70.8 51.6
Benchmark: MNumGLUESub
MathOctopus 7B 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 46.9 37.1
+ PPO LoRA 34.3 41.1 40.7 45.6 49.3 39.5 34.3 46.7 36.3 51.6 41.9
+ DPO Iter1 32.6 41.4 42.0 42.7 45.0 42.2 40.7 47.8 40.9 47.8 42.3
+ DPO Iter2 35.6 39.7 42.2 45.6 50.7 44.1 42.6 49.3 42.9 49.7 44.2
+ DPO Iter3 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 55.9 50.4
MetaMathOctopus 7B 34.7 41.4 37.9 47.8 54.2 45 43.1 52.5 45.0 60.8 46.3
+ PPO LoRA 42.4 46.7 45.4 55.2 58.9 45.0 41.6 55.7 44.6 64.6 50.0
+ DPO Iter1 37.7 46.3 43.5 54.8 58.4 50.5 52.2 60.1 51.6 62.1 51.7
+ DPO Iter2 40.7 46.9 45.0 53.5 58.4 50.8 51.4 58.4 49.9 64.6 52.0
+ DPO Iter3 40.7 46.0 45.0 58.2 59.3 53.1 51.4 57.4 52.0 66.1 52.9
Table 7: Model Performances on three benchmarks. We report the results of PPO LoRA and each round in Iterative DPO. “Avg” represents the average performance on ten languages